OpenStudy (anonymous):

find the point on the curve y=2x^2-3x+1 where the tangent has the gradient 1.

OpenStudy (anonymous):

Find first derivative and set it equal to 1.

OpenStudy (anonymous):

6x^2-x=1?

OpenStudy (anonymous):

*6x^2-3=1?

OpenStudy (anonymous):

Not sure where you're getting the 6..

OpenStudy (anonymous):

Is it actually y = 2x^3 + ... ?

OpenStudy (anonymous):

oh no wait, i meant 4... i read it wrong

OpenStudy (anonymous):

Ok, so your first derivative is y' = 4x-3, right?

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

(I don't know what "yh" means) If so, then find x when 4x-3 = 1.

OpenStudy (anonymous):

it means yeah as in yes ok

OpenStudy (anonymous):

x=13/4?

OpenStudy (anonymous):

Try that again.

OpenStudy (anonymous):

sorry, its 1 isnt it

OpenStudy (anonymous):

i really dont get how i got the previous answer aha

OpenStudy (anonymous):

and y=0?

OpenStudy (anonymous):

Yep, that's it. The point is (1,0)

OpenStudy (anonymous):

ok thanks for the help:)